A 555-mile, 5-hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105 mph. For the remainder of the trip, the average speed was 115 mph. How long did the plane fly at each speed?
Answer by stanbon(75887)

A 555-mile, 5-hour plane trip was flown at two speeds. For the first part of the trip, the average speed was 105 mph. For the remainder of the trip, the average speed was 115 mph. How long did the plane fly at each speed?
--------------------------------------
1st part DATA:
rate = 105 mph ; time = x hrs ; distance = rt = 105x miles
--------------------------------------
2nd part DATA:
rate = 115 mph ; time = 5-x hrs ; distance = rt = 115(5-x) miles
--------------------------------------
Equation:
distance + distance = 555 miles
105x + 115*5 - 115x = 555
-10x = 555-575
x = 2 hrs (time flown at 105 mph)
5-x = 3 hrs (time flown at 115 mph)
==========================================
Cheers,
Stan H.
Reply to stanbon@comcast.net